# GradMix: Gradient-based Selective Mixup for Robust Data Augmentation in Class-Incremental Learning

### v1

@@ 4.2 Empirical Analysis
Figure 2: Comparison of task-wise accuracy over all tasks and class-wise accuracy on the last task between ER and Mixup on the MNIST and CIFAR-10 datasets.
For MNIST, Mixup degrades the accuracy both in task-wise and class-wise. For CIFAR-10, Mixup helps improve task-wise accuracy, but it reduces the accuracy for some previous classes (4th and 5th) at the last task. The accuracy of the current classes (8th and 9th) is not affected by Mixup due to the sufficient amount of current task data compared to the limited buffer data from previous tasks. We find that 51.4% and 47.2% of mixed samples update the model with gradients that satisfy the sufficient condition for greater catastrophic forgetting (in Theorem 1) on the MNIST and CIFAR-10 datasets, respectively. Although Mixup sometimes improves the accuracy, we believe that the gain primarily comes from augmenting the limited buffer data, rather than from mitigating catastrophic forgetting. The results for the Fashion-MNIST, CIFAR-100, and Tiny ImageNet datasets are similar and shown in Sec. C.2. Based on our empirical analysis, we believe that greater catastrophic forgetting may caused by the unpredictable gradients of randomly mixed samples, and the accuracy can be further improved by using a selective mixup. In the next section, we theoretically derive the gradients of mixed samples and propose a selective mixup algorithm to mitigate catastrophic forgetting using these gradients.
5 GradMix
5.1 Problem Formulation
@@
$\underset{d_{j}\in T_{l}\cup\mathcal{M}}{\arg\min}\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\tilde{\ell}(f_{\theta},G_{y}),$ (3)
where the updated loss after training with the mixed sample $\tilde{d}_{ij}$ is approximated as ~$\tilde{\ell}(f_{\theta},G_{y})=\ell(f_{\theta}^{l-1},G_{y})-\eta\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij})$ for a group of data $G_{y}$ by using Eq. 1. Since $\ell(f_{\theta}^{l-1},G_{y})$ and $\eta$ are constant values, we can ignore them and change the argmin to an argmax:
$\underset{d_{j}\in T_{l}\cup\mathcal{M}}{\arg\max}\biggl{(}\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})\biggr{)}^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij}),$ (4)
@@
5.2 Gradients of Mixed Samples
We first compute the approximated gradient of the original sample by employing the last layer approximation DBLP:conf/icml/KatharopoulosF18 ; DBLP:conf/iclr/AshZK0A20 ; DBLP:conf/icml/MirzasoleimanBL20 that only considers the last layer gradients, which is known to be reasonable and efficient DBLP:conf/aaai/KillamsettySRI21 ; DBLP:conf/icml/KillamsettySRDI21 ; DBLP:conf/aaai/KimHW24 . The following Lemma derives the approximated gradient vector of the original sample.
@@ 5.2 Gradients of Mixed Samples
$g=(\nabla_{b}\ell(f_{\theta}^{l-1},d),\nabla_{w}\ell(f_{\theta}^{l-1},d))^{\top}=(\hat{y}-y,(\hat{y}-y)^{\top}X)^{\top},$ (5)
@@
The proof is in Sec. A.2. We next derive the approximated gradient vector of a mixed sample. Similar to the works that theoretically analyze Mixup DBLP:conf/iclr/ZhangDKG021 ; DBLP:conf/icml/Yao0LZL0F22 ; DBLP:conf/icml/ZhangDK022 , suppose we obtain a linear model from the training data. Specifically, the model follows the Fisher’s linear discriminant analysis 10.5555/59551 ; 0ceee346-049e-3934-a519-92e72b194bcf ; 10.1214/20-AOS2012 where the optimal classification rule is linear. Under these assumptions, we present the following theorem that derives the approximated gradient vector of a mixed sample.
@@
$\tilde{g}_{ij}$ $=(\nabla_{b}\ell(f_{\theta}^{l-1},\tilde{d}_{ij}),\nabla_{w}\ell(f_{\theta}^{l-1},\tilde{d}_{ij}))^{\top}$ (6)
@@
The proof is in Sec. A.2. If the value of $\lambda$ is 0 or 1, $\tilde{g}_{ij}$ corresponds to the original gradient of $g_{j}$ or $g_{i}$, respectively. However, if the model is nonlinear, and the value of $\lambda$ is between 0 and 1, there may be an approximation error between the approximated gradient and the true gradient for the mixed sample. Since we use the inner product of the average gradient of buffer data and the gradient of mixed training sample as the objective function in Eq. 4, we empirically compare both the inner product values and the resulting optimal sample pairs when using the approximated gradient and the true gradient for the mixed sample. In practice, the average error in the inner product values is quite small, resulting in high precision and recall for the optimal sample pairs as shown in Sec. C.3.
5.3 Algorithm
We now design the GradMix algorithm to solve the data selection optimization problem. By substituting Eq. 5 and Eq. 6 into Eq. 4 and using the last layer approximation, we can efficiently compute the inner products of the average gradient of buffer data and the gradients of all mixed samples to find the optimal mixing sample. For better efficiency, we group samples that have similar gradients and reformulate the optimization problem from data selection to group selection by replacing the gradient of a mixed sample in Eq. 4 with the average gradient of mixed samples from arbitrary two groups. By leveraging the observation that training samples within the same class share similar gradients as described in Sec. C.4, we group samples by class and propose a class-based selective mixup criterion. If the inner product of the average gradients of buffer data and mixed samples from two classes is negative, it implies that training with the mixed samples increases the average loss of previous tasks and worsens catastrophic forgetting. In this case, we define the class pair as a negative class pair and replace the mixed samples with those generated from the optimal class pair that maximizes the inner product. Conversely, if the inner product is positive, indicating that the average loss of previous tasks does not increase, and catastrophic forgetting does not occur, we retain the mixed samples as part of the training data.
We describe the overall process of GradMix in Algorithm 1 (see Sec. B). At each epoch during training, we compute the average gradient of buffer data (denoted as $\overline{g}_{buf}$) using Eq. 5 (Step 3). Next, we set a mixing ratio $\lambda$ and compute the average gradients of mixed samples from all class pairs (denoted as $\tilde{g}_{y_{i}y_{j}}$) using Eq. 6 (Steps 4–7). By computing the inner products of $\overline{g}_{buf}$ and $\tilde{g}_{y_{i}y_{j}}$ for all class pairs, we get the set of negative class pairs (Step 8) and the set of optimal class pairs for each class (Step 9). The selective mixup criterion is determined using the two sets once per epoch and applied to all batches to ensure computational efficiency. For each batch consisting of both current task data and buffer data (Steps 11–13), we first randomly shuffle the batch to generate sample pairs to be mixed and set a mixing ratio $\lambda$ following the configuration in Mixup DBLP:conf/iclr/ZhangCDL18 (Steps 14–15). However, if the class pairs of the sample pairs belong to the negative set, we replace them with sample pairs from the optimal class pairs (Steps 16–19). After this replacement process, we form a new batch consisting of mixed samples that satisfy our selective mixup criterion (Step 22) and update the model (Step 23). Lastly, we store a random subset of the current task data to the buffer before moving onto the next task (Steps 26–27). We provide an analysis of the computational complexity and runtime of GradMix algorithm in Sec. C.5, which shows that GradMix is scalable by only computing the last layer gradients of a small exemplar set for each class.
6 Experiments
Metrics
Table 1: For each of the five datasets, we show the total training set size, the number of features, the number of classes, and the number of tasks.
Dataset Size #Ftrs #Cls #Tsks
MNIST 50K 28$\times$28 10 5
FMNIST 60K 28$\times$28 10 5
CIFAR-10 50K 3$\times$32$\times$32 10 5
CIFAR-100 50K 3$\times$32$\times$32 100 10
Tiny ImageNet 100K 3$\times$64$\times$64 200 10
Datasets
We use a total of five datasets for the experiments as shown in Table 1: MNIST DBLP:journals/pieee/LeCunBBH98 , Fashion-MNIST (FMNIST) DBLP:journals/corr/abs-1708-07747 , CIFAR-10 Krizhevsky2009LearningML , CIFAR-100 Krizhevsky2009LearningML , and Tiny ImageNet DBLP:journals/ijcv/RussakovskyDSKS15 . To setup experiments for class-incremental learning, we divide the datasets into sequences of tasks consisting of equal numbers of disjoint classes and assume that the task boundaries are available DBLP:journals/corr/abs-1904-07734 ; DBLP:conf/nips/MirzadehFPG20 .
Models
For MNIST and FMNIST, we train a two-layer MLP each with 256 neurons using the SGD optimizer with an initial learning rate of 0.01 for 20 epochs. For CIFAR-10, CIFAR-100, and Tiny ImageNet, we train a ResNet-18 DBLP:conf/cvpr/HeZRS16 ; DBLP:conf/eccv/HeZRS16 model using the SGD optimizer with an initial learning rate of 0.1 for 50, 250, and 100 epochs, respectively. During training, we reduce the learning rate by one-tenth at 100, 150, and 200 epochs for CIFAR-100, and at 30, 60, and 90 epochs for Tiny ImageNet. The last layer of the model is shared for all the tasks for a single-head evaluation DBLP:journals/corr/abs-1805-09733 ; DBLP:conf/eccv/ChaudhryDAT18 . We train all models using cross-entropy loss.
Baselines
Following the experimental setups of DBLP:conf/nips/KumariW0B22 ; DBLP:conf/nips/ZhangPFBLJ22 , we use ER DBLP:journals/corr/abs-1902-10486 , MIR DBLP:conf/nips/AljundiBTCCLP19 , GSS DBLP:conf/nips/AljundiLGB19 , and DER DBLP:conf/nips/BuzzegaBPAC20 as basis experience replay methods for class-incremental learning and apply data augmentation techniques to them. The experiments also can be extended to include other experience replay methods. We compare GradMix with three types of data augmentation baselines:
  • Mixup-based methods: MixupDBLP:conf/iclr/ZhangCDL18 generates synthetic training data through linear interpolations of pairs of input features and their corresponding labels; Manifold MixupDBLP:conf/icml/VermaLBNMLB19 performs linear interpolations of hidden layer representations and corresponding labels of training data; CutMixDBLP:conf/iccv/YunHCOYC19 generates new training data by randomly cutting and pasting patches between input images, while adjusting the corresponding labels based on the proportion of each patch.
  • Imbalance-aware mixup methods: RemixDBLP:conf/eccv/ChouCPWJ20 mixes features like Mixup, but assigns labels by providing higher weights to the minority class; Balanced-MixUpDBLP:conf/miccai/GaldranCB21 performs both regular and balanced sampling when selecting pairs of training data for Mixup.
  • Policy-based method: RandAugmentDBLP:conf/nips/CubukZS020 randomly applies different image transformations to training data with adjustable magnitude.
Table 2: Average accuracy results on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use ER as a basis method of class-incremental learning and compare GradMix with three types of data augmentation baselines: (1) mixup-based methods: Mixup, Manifold Mixup, and CutMix; (2) imbalance-aware mixup methods: Remix and Balanced-MixUp; and (3) a policy-based method: RandAugment.
Type Method MNIST FMNIST CIFAR-10 CIFAR-100 Tiny ImageNet
Experience Replay ER 0.896$\pm$0.002 0.756$\pm$0.004 0.613$\pm$0.009 0.491$\pm$0.019 0.385$\pm$0.009
Mixup-based + Mixup 0.866$\pm$0.005 0.766$\pm$0.009 0.643$\pm$0.008 0.507$\pm$0.010 0.371$\pm$0.008
+ Manifold Mixup 0.896$\pm$0.004 0.774$\pm$0.004 0.620$\pm$0.011 0.521$\pm$0.008 0.386$\pm$0.009
+ CutMix 0.878$\pm$0.005 0.789$\pm$0.004 0.516$\pm$0.009 0.404$\pm$0.009 0.344$\pm$0.010
Imbalance-aware + Remix 0.905$\pm$0.004 0.789$\pm$0.005 0.639$\pm$0.014 0.528$\pm$0.007 0.391$\pm$0.008
+ Balanced-MixUp 0.883$\pm$0.006 0.767$\pm$0.006 0.630$\pm$0.014 0.516$\pm$0.009 0.379$\pm$0.008
Policy-based + RandAugment 0.765$\pm$0.008 0.640$\pm$0.008 0.636$\pm$0.006 0.484$\pm$0.007 0.368$\pm$0.007
Selective Mixup + GradMix (ours) 0.918$\pm$0.004 0.802$\pm$0.009 0.667$\pm$0.025 0.546$\pm$0.008 0.405$\pm$0.006
Parameters
We store 32 samples per class for buffering and set the batch size to 64 for all experiments. For the hyperparameter $\alpha$ used in the beta distribution, we follow the setting of Mixup DBLP:conf/iclr/ZhangCDL18 . Since the Mixup paper does not specify $\alpha$ values for the MNIST and FMNIST datasets, we set $\alpha$ to 1.0, which corresponds to a uniform distribution. In practice, the performance of GradMix is not sensitive to the value of $\alpha$ as shown in Sec. C.6. See Sec. C.1 for more experimental details.
6.1 Model Performance Results
We compare GradMix against other data augmentation baselines on the five datasets with respect to average accuracy as shown in Table 2 and Sec. C.7. Overall, GradMix consistently achieves the best accuracy results compared to the baselines for all the datasets by considering catastrophic forgetting while augmenting data. In contrast, since existing data augmentation techniques are designed for static and balanced data, they have limitations in class-incremental learning setups with dynamic and imbalanced data. In addition, GradMix helps further improvement of advanced experience replay methods, including MIR, GSS, and DER, as shown in Table 3.
Among the three types of baselines, the two imbalance-aware mixup methods, Remix and Balanced-MixUp, achieve slightly better results than other types of baselines as they address the data imbalance between previous and current tasks. However, the random selection of sample pairs for mixup leads to greater catastrophic forgetting of previous tasks and limits the improvement in average accuracy. In comparison, GradMix performs gradient-based selective mixup to minimize catastrophic forgetting of previous tasks, which leads to significantly better model accuracy results for all datasets. See a more detailed analysis of the baselines in Sec. C.8. We present sequential accuracy results for each task as shown in Sec. C.9 where GradMix also shows better accuracy performance than other baselines at each task. We also show that GradMix is effective on large-scale datasets (e.g., ImageNet-1K DBLP:conf/cvpr/DengDSLL009 ) and variations in backbone models (e.g., ViT DBLP:conf/iclr/DosovitskiyB0WZ21 ) in Sec. C.10 and Sec. C.11, respectively.
Table 3: Average accuracy results on the CIFAR-10, CIFAR-100, and Tiny ImageNet datasets. We use MIR, GSS, and DER as basis methods of class-incremental learning and integrate GradMix.
Method CIFAR-10 CIFAR-100 Tiny ImageNet
MIR 0.620$\pm$0.009 0.493$\pm$0.021 0.386$\pm$0.009
+ GradMix 0.669$\pm$0.013 0.546$\pm$0.007 0.411$\pm$0.009
GSS 0.627$\pm$0.009 0.512$\pm$0.013 0.397$\pm$0.008
+ GradMix 0.674$\pm$0.011 0.563$\pm$0.007 0.418$\pm$0.006
DER 0.678$\pm$0.010 0.560$\pm$0.009 0.422$\pm$0.008
+ GradMix 0.714$\pm$0.009 0.604$\pm$0.006 0.452$\pm$0.006
(a) MNIST.
(b) CIFAR-10.
Figure 3: Selective mixup results of GradMix on the last task for MNIST and CIFAR-10 datasets.
Table 4: Ablation study on the MNIST, FMNIST, CIFAR-10, CIFAR-100, and Tiny ImageNet datasets with two ablation scenarios when randomly selected sample pairs worsen catastrophic forgetting in Algorithm 1 (Steps 17–19): (1) using original samples without mixup; and (2) retaining random mixup. We use ER as a basis method.
Method MNIST FMNIST CIFAR-10 CIFAR-100 Tiny ImageNet
Original 0.904$\pm$0.005 0.773$\pm$0.005 0.621$\pm$0.008 0.518$\pm$0.010 0.383$\pm$0.010
Random 0.905$\pm$0.004 0.789$\pm$0.005 0.643$\pm$0.008 0.528$\pm$0.007 0.391$\pm$0.008
GradMix (ours) 0.918$\pm$0.004 0.802$\pm$0.009 0.667$\pm$0.025 0.546$\pm$0.008 0.405$\pm$0.006
Table 5: Average accuracy results on the FMNIST, CIFAR-10, and CIFAR-100 datasets with respect to the buffer size per class. We use ER as a basis method and compare GradMix with three types of data augmentation baselines. The buffer size is varied from 32 to 16 and 64 samples per class.
Method FMNIST CIFAR-10 CIFAR-100
$N$=16 $N$=64 $N$=16 $N$=64 $N$=16 $N$=64
ER 0.716$\pm$0.006 0.792$\pm$0.004 0.546$\pm$0.022 0.682$\pm$0.005 0.412$\pm$0.021 0.555$\pm$0.021
+ Mixup 0.733$\pm$0.006 0.811$\pm$0.005 0.569$\pm$0.022 0.718$\pm$0.010 0.413$\pm$0.014 0.587$\pm$0.016
+ Manifold Mixup 0.739$\pm$0.004 0.818$\pm$0.003 0.563$\pm$0.012 0.688$\pm$0.011 0.419$\pm$0.013 0.603$\pm$0.011
+ CutMix 0.743$\pm$0.011 0.830$\pm$0.003 0.475$\pm$0.009 0.602$\pm$0.016 0.347$\pm$0.007 0.505$\pm$0.006
+ Remix 0.751$\pm$0.006 0.825$\pm$0.005 0.565$\pm$0.017 0.711$\pm$0.011 0.423$\pm$0.010 0.607$\pm$0.012
+ Balanced-MixUp 0.724$\pm$0.005 0.818$\pm$0.005 0.560$\pm$0.028 0.696$\pm$0.006 0.426$\pm$0.016 0.597$\pm$0.016
+ RandAugment 0.600$\pm$0.008 0.696$\pm$0.005 0.584$\pm$0.008 0.703$\pm$0.010 0.415$\pm$0.013 0.560$\pm$0.009
+ GradMix (ours) 0.764$\pm$0.009 0.837$\pm$0.007 0.588$\pm$0.021 0.739$\pm$0.009 0.437$\pm$0.009 0.620$\pm$0.011
6.2 Selective Mixup Analysis
We next analyze how GradMix selectively mixes different pairs of classes. We provide selective mixup results on the last task of the MNIST and CIFAR-10 datasets as shown in Fig. 3, and the results for the FMNIST, CIFAR-100, and Tiny ImageNet datasets are shown in Sec. C.12. Since the selective mixup results for other tasks are similar, we show the results for the last task without loss of generality. As the selective mixup criterion changes each epoch based on gradients, we sum the total number of mixup operations for each class pair combination, denoted as $(C_{i},C_{j})$, during training and present the results as a heatmap.
6.3 Ablation Study
6.4 Varying the Buffer Size
We compare the performance of GradMix with the baselines when varying the number of buffer samples per class, denoted as $N$, from 32 to 16 and 64 as shown in Table 5. We observe that GradMix shows the best performance for all datasets regardless of the buffer size. In addition, as we store more buffer samples and use them for training, the average accuracy of all the baselines and GradMix increases. The results for the MNIST and Tiny ImageNet datasets are similar and shown in Sec. C.13.
7 Conclusion
Appendix A Appendix – Theory
A.1 Theoretical Analysis of Applying Mixup in Class-Incremental Learning
Theorem.
Proof.
$L_{\text{ER}}$ $=\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\tilde{\ell}(f_{\theta},G_{y})$ (7)
$L_{\text{Mixup}}=\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\ell(f_{\theta}^{l-1},G_{y})-\eta\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij}).$ (8)
By subtracting Eq. 8 from Eq. 7,
$L_{\text{ER}}-L_{\text{Mixup}}$ $=-\eta\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\Bigl{(}\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},d_{i})-\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij})\Bigr{)}$ (9)
$L_{\text{ER}}-L_{\text{Mixup}}<0.$
A.2 Theoretical Analysis of the Gradients of Mixed Samples
Lemma.
Proof.
$\hat{y}=\sigma(z)=\sigma(w\cdot X+b).$
$\frac{\partial\ell}{\partial z_{j}}=-\frac{\partial}{\partial z_{j}}\sum_{i=1}^{c}y_{i}\cdot\log(\hat{y}_{i})=-\sum_{i=1}^{c}y_{i}\cdot\frac{\partial}{\partial z_{j}}\log(\hat{y}_{i}).$ (10)
$\frac{\partial}{\partial z_{j}}\log(\hat{y}_{i})$ $=\frac{\partial}{\partial z_{j}}\log\biggl{(}\frac{e^{z_{i}}}{\sum_{n=1}^{c}e^{z_{n}}}\biggr{)}$ (11)
$=\mathbbm{1}\{i=j\}-\frac{e^{z_{j}}}{\sum_{n=1}^{c}e^{z_{n}}}$
$=\mathbbm{1}\{i=j\}-\hat{y}_{j},$
$\frac{\partial\ell}{\partial z_{j}}$ $=-\sum_{i=1}^{c}y_{i}\cdot(\mathbbm{1}\{i=j\}-\hat{y}_{j})$ (12)
$=\sum_{i=1}^{c}y_{i}\cdot\hat{y}_{j}-\sum_{i=1}^{c}y_{i}\cdot\mathbbm{1}\{i=j\}$
$\frac{\partial\ell}{\partial z_{j}}=\hat{y}_{j}-y_{j}.$
In the vector notation, we get:
$\frac{\partial\ell}{\partial z}=\hat{y}-y.$
$g=(\nabla_{b}\ell(f_{\theta}^{l-1},d),\nabla_{w}\ell(f_{\theta}^{l-1},d))^{\top}=(\hat{y}-y,(\hat{y}-y)^{\top}X)^{\top}.$ (13)
Theorem.
Proof.
$\tilde{g}_{ij}$ $=(\hat{y}_{ij}-\tilde{y}_{ij},(\hat{y}_{ij}-\tilde{y}_{ij})^{\top}X_{ij})^{\top}$ (14)
$\hat{y}=\sigma(f_{\theta}^{l-1}(x))=\sigma(\theta^{\top}x).$
Then, the logits for $x_{i}$ and $x_{j}$ become:
$z_{ij}$ $=\theta^{\top}\tilde{x}_{ij}$
$=\theta^{\top}(\lambda x_{i}+(1-\lambda)x_{j})$
$=\lambda\theta^{\top}x_{i}+(1-\lambda)\theta^{\top}x_{j}$
$\hat{y}_{ij}$ $=\sigma(z_{ij})$ (15)
$X_{ij}=\lambda X_{i}+(1-\lambda)X_{j}.$ (16)
Appendix B Appendix – Algorithm
Continuing from Sec. 5.3, we provide the full algorithm of GradMix.

### v2

@@ 4.1. Theoretical Analysis
Let $\tilde{d}_{ij}$ be a mixed sample by mixing an original training sample $d_{i}$ and a randomly selected sample $d_{j}$. If the gradient of the mixed sample satisfies the following condition:
$\sum_{y\in\mathbb{Y}_{p}}\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\left(\nabla_{\theta}\ell(f_{\theta}^{l-1},d_{i})-\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij})\right)>0,$
then training with the mixed sample leads to a higher average loss for previous tasks and worsens catastrophic forgetting compared to training with the original sample.
@@ 4.2. Empirical Analysis
Figure 2. Comparison of task-wise accuracy over all tasks and class-wise accuracy on the last task between ER and Mixup on the MNIST and CIFAR-10 datasets.
For MNIST, Mixup degrades the accuracy both in task-wise and class-wise. For CIFAR-10, Mixup helps improve task-wise accuracy, but it reduces the accuracy for some previous classes (4th and 5th) at the last task. The accuracy of the current classes (8th and 9th) is not affected by Mixup due to the sufficient amount of current task data compared to the limited buffer data from previous tasks. We find that 51.4% and 47.2% of mixed samples update the model with gradients that satisfy the sufficient condition for greater catastrophic forgetting (in Theorem 1) on the MNIST and CIFAR-10 datasets, respectively. Although Mixup sometimes improves the accuracy, we believe that the gain primarily comes from augmenting the limited buffer data, rather than from mitigating catastrophic forgetting. The results for the Fashion-MNIST, CIFAR-100, and Tiny ImageNet datasets are similar and shown in Sec. B.2. Based on our empirical analysis, we believe that greater catastrophic forgetting may be caused by the unpredictable gradients of randomly mixed samples, and the accuracy can be further improved by using a selective mixup. In the next section, we theoretically derive the gradients of mixed samples and propose a selective mixup algorithm to mitigate catastrophic forgetting using these gradients.
5. GradMix
5.1. Problem Formulation
@@
(3) $\underset{d_{j}\in T_{l}\cup\mathcal{M}}{\arg\min}\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\tilde{\ell}(f_{\theta},G_{y}),$
where the updated loss after training with the mixed sample $\tilde{d}_{ij}$ is approximated as
(4) $\tilde{\ell}(f_{\theta},G_{y})=\ell(f_{\theta}^{l-1},G_{y})-\eta\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij})$
for a group of data $G_{y}$ by using Eq. 1. Since $\ell(f_{\theta}^{l-1},G_{y})$ and $\eta$ are constant values, we can ignore them and change the argmin to an argmax:
(5) $\underset{d_{j}\in T_{l}\cup\mathcal{M}}{\arg\max}\biggl(\frac{1}{|\mathbb{Y}_{p}|}\sum_{y\in\mathbb{Y}_{p}}\nabla_{\theta}\ell(f_{\theta}^{l-1},G_{y})\biggr)^{\top}\nabla_{\theta}\ell(f_{\theta}^{l-1},\tilde{d}_{ij}),$
@@
5.2. Gradients of Mixed Samples
We first compute the approximated gradient of the original sample by employing the last layer approximation (Katharopoulos and Fleuret, 2018; Ash et al., 2020; Mirzasoleiman et al., 2020) that only considers the last layer gradients, which is known to be reasonable and efficient (Killamsetty et al., 2021a, b; Kim et al., 2024b). The following Lemma derives the approximated gradient vector of the original sample.
@@ 5.2. Gradients of Mixed Samples
(6) $g=(\nabla_{b}\ell(f_{\theta}^{l-1},d),\nabla_{w}\ell(f_{\theta}^{l-1},d))^{\top}=(\hat{y}-y,(\hat{y}-y)^{\top}X)^{\top},$
@@
The proof is in Sec. A.2. We next derive the approximated gradient vector of a mixed sample. Similar to the works that theoretically analyze Mixup (Zhang et al., 2021; Yao et al., 2022; Zhang et al., 2022a), suppose we obtain a linear model from the training data. Specifically, the model follows the Fisher’s linear discriminant analysis (Johnson and Wichern, 1988; Cai and Zhang, 2019, 2021) where the optimal classification rule is linear. Under these assumptions, we present the following theorem that derives the approximated gradient vector of a mixed sample.
@@
(7) $\tilde{g}_{ij}$ $=(\nabla_{b}\ell(f_{\theta}^{l-1},\tilde{d}_{ij}),\nabla_{w}\ell(f_{\theta}^{l-1},\tilde{d}_{ij}))^{\top}$
$=\Bigl(\lambda(\hat{y}_{ij}-y_{i})+(1-\lambda)(\hat{y}_{ij}-y_{j}),$
$\hskip 13.00005pt(\lambda(\hat{y}_{ij}-y_{i})+(1-\lambda)(\hat{y}_{ij}-y_{j}))^{\top}X_{ij}\Bigr)^{\top}$
@@
The proof is in Sec. A.2. If the value of $\lambda$ is 0 or 1, $\tilde{g}_{ij}$ corresponds to the original gradient of $g_{j}$ or $g_{i}$, respectively. However, if the model is nonlinear, and the value of $\lambda$ is between 0 and 1, there may be an approximation error between the approximated gradient and the true gradient for the mixed sample. Since we use the inner product of the average gradient of buffer data and the gradient of mixed training sample as the objective function in Eq. 5, we empirically compare both the inner product values and the resulting optimal sample pairs when using the approximated gradient and the true gradient for the mixed sample. In practice, the average error in the inner product values is quite small, resulting in high precision and recall for the optimal sample pairs as shown in Sec. B.3.